I am having trouble with this one can someone help Thank you so much!
6x – 0.37 = 5(x – 0.28)

Answer by bonster(299)

6x – 0.37 = 5(x – 0.28)
6x – 0.37 = 5x-1.40 <--add .37 to both sides
(6x – 0.37)+0.37 = (5x-1.40)+0.37
6x=5x-1.03 <--subtract 5x from both sides
(6x)-5x=(5x-1.03)-5x
x=-1.03